OpenStudy (anonymous):

STATISTICS HELP PLEASE I really need some help on this one I have been on it for hours ans still cant get it right.... Below are 36 sorted ages of an acting award winner. Find P80 20, 22, 22, 24, 25, 28, 28, 29,29,30,34,35,35,36,37,39,44,48,49,51,51,51,54,56,57,58,63,66,67,72,72,74,76,77,79,80 It keeps giving me the equation L= k/100 * n and I have tried it every way possible please help.

Directrix (directrix):

Please post the entire problem. What is meant by >>Find P80 Also, please define the variaables L and k and give the significance or origin of the equation L= k/100 * n. Thanks.

Directrix (directrix):

Is this the question (see attached) except that we are looking for the age at the 80th percentile or something to do with the 80th percentile?

Directrix (directrix):

If the question is to find the percentile of the award age of 80, then there are 35 scores preceding 80. So, 35/36 = k/100 ----> Cross multiply and solve for k Add a % symbol. Note: I am making all sorts of assumptions about the problem. Therefore, don't assume that my partial solution is correct until you determine if my assumptions mesh with those of the given problem in your text. Note2: L= k/100 * n -> I rewrote this as: L/n = k/100.